f the magnitude of the resultant force is to be 9kN directed along the




positive x axis, determine the magnitude of force T acting on the eyebolt




and its angle delta.

1
Expert's answer
2021-12-02T10:05:20-0500

"T=\sqrt{{F_1}^2+{F_2}^2-2F_1F_2\cos45°},"

"T=6.57~kN,"


"\delta= \arctan \frac {F_1\sin 45°}{T},"

"\delta= 30.6°."